OpenStudy (anonymous):

find the exact value of cot(13/6), with out using a calculator?

OpenStudy (anonymous):

i am going to guess it is \[\cot(\frac{13\pi}{6})\]

OpenStudy (anonymous):

yes sorry . I have the answer in the back of the book as -1.. just not sure how to get there.

OpenStudy (anonymous):

first note that \(\frac{13\pi}{6}\) is coterminal with \(\frac{13\pi}{6}-2\pi=\frac{\pi}{6}\)

OpenStudy (anonymous):

then find \(\cot(\frac{\pi}{6})=\frac{\cos(\frac{\pi}{6})}{\sin(\frac{\pi}{6})}\)and it is not \(-1\) for sure

OpenStudy (anonymous):

ohh I read the back and its sqrt3!

OpenStudy (anonymous):

that makes more sense

OpenStudy (anonymous):

\[\frac{\frac{\sqrt3}{2}}{\frac{1}{2}}=\sqrt3\]

OpenStudy (anonymous):

thankkkkkkkkkkkkkkkk youuuuuuuuuuu!!!!

OpenStudy (anonymous):

yw